# HiBO: Hierarchical Bayesian Optimization via Adaptive Search Space Partitioning

### v4

@@ 6 Evaluation
In this section, we present our experiments including: 1) evaluation on synthetic benchmarks, to investigate the basic effectiveness of HiBO and compare it with related works; 2) a real-world case study of automatic DBMS configuration tuning to demonstrate its performance and practical value on high-dimensional, real-world problems; and 3) a visual analysis of the distribution of resulting partition-score-weighted acquisition function values to validate the behaviours of HiBO. The choice of real-world benchmark is discussed in Section 6.2.1.
@@ 6.1.1 Experiment Setup
To evaluate HiBO on synthetic benchmarks, we compare it with several related approaches, including vanilla GP-based BO, TuRBO (Eriksson et al., 2019), LA-MCTS (Wang et al., 2020), MCTS-VS (Song et al., 2022) and CMA-ES (Lozano et al., 2006). TuRBO is selected for validating the effects of the extra space partitioning information. We compare with LA-MCTS, the current SOTA in space-partitioning-based high-dimensional BO, to validate our design’s efficiency, while excluding prior methods like HOO and LA-NAS for simplicity. We also include MCTS-VS, which uses MCTS for variable selection, and CMA-ES as a baseline evolutionary algorithm. Besides, for evaluating the effectiveness of HiBO on guiding other local optimizer, we introduce HiBO-GP, where HiBO coordinates with the vanilla GP-based BO as the local optimizer.
@@
We evaluate these algorithms on six high-dimensional synthetic benchmarks, including 200-dimensional synthetic functions Rastrigin and Levy with 200 dimensions. To evaluate how well the algorithms focus on relevant dimensions in sparse tasks, we evaluate Branin2-500D, Hartmann6-300D, Rastrigin20-200D, and Levy20-200D111’Benchmark(x)-(y)D’ indicates $x$ effective dimensions out of $y$, with the remaining $(y-x)$ being dummy dimensions.. We run 1000 iterations for Branin2-500D and 500 iterations for the others, starting with 40 and 20 iterations of random sampling, respectively. The additional configuration details for these algorithms are shown in Appendix A.1.
6.1.2 Results
As observed from results on all these benchmarks (Figure 3), HiBO consistently achieves the best result across all benchmarks. Its superior performance compared to TuRBO demonstrates the advantage of incorporating global space partitioning information into the local model’s acquisition strategy. HiBO also outperforms LA-MCTS on most benchmarks, indicating the effectiveness of navigating the local optimizer with partitioning information rather than limiting the sampling scope to the most promising partition. The performance gap with other algorithms widens on sparse benchmarks and becomes greatest on Hartmann6-300D, where HiBO achieves over two magnitudes lower regrets than others. This indicates HiBO’s strong capacity to identify the most promising regions in high-dimensional tasks. While HiBO shows similar performance to LA-MCTS on benchmarks where all dimensions are relevant, like Levy-200D, it finally finds higher-quality solutions. This suggests that HiBO may require additional iterations to adjust its partitioning strategy to cover all dimensions in dense benchmarks but can still outperform in finding optimal solutions afterwards. On the other hand, HiBO-BO is also observed to perform better than GP-BO to different extents on varying benchmarks. It can even achieve the second best and surpasses LA-MCTS in Branin2-500D task. But the improvement of HiBO-GP over GP-BO is not as stable as HiBO over TuRBO. We attribute this to differences in the local optimizer and the further guidance from HiBO can more likely produce consistent improvements over better local optimizers.
@@ 6.2.1 Evaluation Background
Besides evaluation on synthetic benchmarks, we evaluate the practical value of HiBO on a more complex real-world application, specifically automatic DBMS configuration tuning. The performance of DBMS highly depends on the settings of their configuration knobs, which control various aspects of their behaviours such as memory allocation, I/O behaviors, and query optimization. However, default configurations are typically suboptimal and require further tuning for optimal performance (Zhao et al., 2023). We select DBMS configuration tuning as a critical evaluation scenario due to its significant practical importance and inherent complexities, especially for the following considerations:
  • (High-dimensionality) DBMS configuration tuning involves a high-dimensional search space with over 100 configuration knobs;
  • (Knob Correlation) Many of these configuration parameters are interdependent, where certain knobs are correlated and can lead to combined effects on the performance;
  • (Noisy Observations) Measurements of performance data in real-world environments are much noiser compared to cases in synthetic benchmarks.
These factors make manual and heuristic tuning impractical (Huang et al., 2019; Zhao et al., 2023), making DBMS configuration tuning an ideal target for evaluating practical values of algorithms in real-world applications. In comparison, though not limited to simple mathematical synthetic functions, commonly-seen “real-world" benchmarks in literature are still large based on simulation (Jones, 2008; Wang et al., 2017; Eriksson and Jankowiak, 2021). For example, Mopta08 (Jones, 2008) originates from real-world problems in vehicle design but is abstracted into a mathematical optimization problem in practice. Considering the page limits, we included the evaluation on two simulated real-world benchmarks in Appendix E, including Mopta08 and Rover Trajectory Planning (Wang et al., 2017).
@@ 6.2.2 Experiment Setup
Our evaluation focuses on the effectiveness of HiBO for DBMS configuration tuning without relying on extra sampling besides its own basic initial sampling. We choose PostgreSQL (version 16.3) as the target DBMS considering its accessibility and general applicability. From OnGres Inc. (2023), we manually selected 110 knobs for tuning, excluding those related to debugging, network connections, and authorization. SysBench (Kopytov, 2004) is applied as a standard OLTP workload in the experiment throughput (measured in events per second, EPS) as the target metric to be maximized. We used a limited sample budget of 50 iterations for simulating scenarios in production environments.
@@ 6.3.1 Results and Analysis
Figure 4 (Left) summarizes the performance improvements of PostgreSQL on SysBench achieved by different algorithms. HiBO shows the highest throughput gains in greatest improvement rate, achieving about 28% of performance improvements over the default performance (200 EPS) than LA-MCTS and TuRBO do in 50 iterations. The comparison between HiBO and TuRBO highlights HiBO’s effectiveness in guiding local models, even in complex and noisy real-world scenarios. LA-MCTS falls behind HiBO and is close to TuRBO, further indicating that simply limiting the sampling scope is insufficient, especially with limited sample budgets. Moreover, without additional pre-sampling, OtterTune* and LlamaTune also fall behind space-partitioning-based approaches.
@@
Besides performance improvements, in real-world scenarios, tuning time and the rate of suggesting failure-causing configurations must be considered towards better practicality. A failure-causing configuration, defined as unsafe (Zhang et al., 2022), can degrade the performance of critical DBMS-dependent processes in the system. Therefore, we introduce the Safety-weighted Performance Improvement-Time Ratio (S-PITR): $S\text{-}PITR=\frac{PI}{TT+NF\cdot PE}$, where $PI$ is performance improvement, $TT$ refers to tuning time, $NF$ represents the number of failed (unsafe) configurations during the search and $PE$ is a the penalty value for failed configurations. This metric quantitatively measures the practicality of tuning algorithms by considering both tuning time and safety.
@@
Figure 4 (Right) presents the S-PITR values for the selected algorithms while tuning PostgreSQL’s configurations towards SysBench. Again, HiBO achieves the highest S-PITR value and outperforms LA-MCTS by over 20%, demonstrating its capability to navigate towards promising regions while accounting for both failure rate and computational cost. In comparison, LA-MCTS even falls behind TuRBO in terms of S-PITR, suggesting that its added complexity does not efficiently translate into practical benefits. Besides, HiBO presents $\sim$25greater S-PITR scores than OtterTune and LlamaTune. These findings highlight HiBO’s adequate ability to strike a balance between performance improvement, safety, and efficiency in complex real-world DBMS tuning scenarios.
@@
We visualize the acquisition function values generated by HiBO and TuRBO during optimization of the Ackley2-200D synthetic function over 100 iterations. Starting with 20 initial samples, we plot the distribution of acquisition values on the two effective dimensions as the optimization progresses. For clarity, the top 1000 acquisition values from 100x100 uniformly distributed samples are shown in each iteration. Figure 5 illustrates these results. The comparison shows that HiBO, with the extra information of space partitioning, more rapidly converges its acquisition distribution towards the optimal point, as indicated by the brightest areas near the optima. In contrast, while also moving toward the optimal region, the convergence speed of TuRBO is way slower and exhibits a less concentrated and noisier final distribution, with significant acquisition values still dispersed far from the optima. This visual analysis further validates the effectiveness of HiBO by integrating space partitioning into the acquisition strategy of the local optimizer.
@@ 7 Conclusion
In this paper, we introduce HiBO, a novel Bayesian Optimization algorithm designed to efficiently handle high-dimensional search spaces via a hierarchical integration of global-level space partitioning and local modeling. HiBO dynamically refines global partitioning information and seamlessly integrates it into local acquisition strategies, thereby accelerating the convergence towards global optima. Our comprehensive evaluation demonstrate its effectiveness and practical value on both synthetic benchmarks and, more importantly, on real-world systems. Additionally, the visual analysis presents a straightforward validation of its effects in navigating on complex optimization landscapes. More discussions on limitations and scalability of HiBO can be found in Appendix F and G respectively.
@@ A.1 Algorithm Configurations for Synthetic Benchmarks
  • GP-BO: Implemented using ‘GaussianProcessRegressor‘ from scikit-learn library (Pedregosa et al., 2011), configured with a Matérn kernel (ν=2.5) (Matérn, 2013) scaled by a constant factor of 1.0, and the noise level is set to 0.1, which translates to an alpha value of 0.01. The acquisition function is set to Expected Improvement (Jones et al., 1998);
@@ A.2 Experiment Setup for DBMS Configuration Tuning
We select PostgreSQL v16.3 (PostgreSQL Global Development Group, 2023) considering its accessibility and generality. PostgreSQL is a powerful, open-source Relational DBMS that has become one of the most popular choices for managing enterprise-level databases (Merit Data Tech, 2023; Percona, 2023) given its reliability and extensibility. Importantly, PostgreSQL exposes up to more than 200 configuration knobs as shown in PGCONF (OnGres Inc., 2023) which makes it an ideal instance for our DBMS configuration tuning experiments. We collected information of all configuration knobs of PostgreSQL in PostgreSQL Co.NF (OnGres Inc., 2023). A total of 110 knobs were manually selected, excluding those related to debugging, network connections, and authorization, as they could interfere with the interaction between the DBMS controller module and the DBMS instance and have negligible impact on DBMS performance.
@@ A.4 SysBench Statement Types
  • Point selection: "SELECT c FROM sbtest%u WHERE id=?"

### v6

@@ 5 Evaluation
In this section, we present our evaluation experiments including: 1) synthetic benchmarks and 2) a real-world case study of automatic DBMS configuration tuning to comprehensive measure the effectiveness and practical value of HiBO on high-dimensional problems; 3) a comparative visual analysis of HiBO’s acquisition function distribution across iterations to validate the behaviours of HiBO; and 4) empirical validation on the scalability of HiBO with respect to increased dimensionality of the search space.
@@ 5.1.1 Experiment Setup
To evaluate HiBO on synthetic benchmarks, we compare it with several related approaches, including vanilla GP-based BO, TuRBO (Eriksson et al., 2019), LA-MCTS (Wang et al., 2020), MCTS-VS (Song et al., 2022) and CMA-ES (Lozano et al., 2006). TuRBO is selected for validating the effects of the extra space partitioning information. We compare with LA-MCTS, the current SOTA in space-partitioning-based high-dimensional BO, to validate our design’s efficiency, while excluding prior methods like HOO and LA-NAS for simplicity. We also include MCTS-VS, which uses MCTS for variable selection. Besides, for evaluating the effectiveness of HiBO on guiding other local optimizer, we introduce HiBO-GP, where the GP-based BO serves as the local optimizer of HiBO.
@@
We evaluate these algorithms on six high-dimensional synthetic benchmarks. To evaluate how well the algorithms focus on relevant dimensions in sparse tasks, we evaluate Branin2-500D, Hartmann6-300D, Rastrigin20-200D, and Levy20-200D222’Benchmark(x)-(y)D’ indicates $x$ effective dimensions out of $y$, with the remaining $(y-x)$ being dummy dimensions.. Here regret is measured for Hartmann6 and Branin2 benchmarks and defined as the difference between the optimal objective value and the algorithm’s achieved value at a given iteration, with lower regret indicating better algorithm performance. We run 1000 iterations for Branin2-500D and 500 iterations for the others, starting with 40 and 20 iterations of random sampling, respectively. The additional configuration details for these algorithms are shown in Appendix A.1.
As observed from results on all these benchmarks (Figure 3), HiBO consistently achieves the best result across all benchmarks. Its superior performance compared to TuRBO demonstrates the advantage of incorporating global space partitioning information into the local model’s acquisition strategy. HiBO also outperforms LA-MCTS on most benchmarks, indicating the effectiveness of navigating the local optimizer with partitioning information rather than limiting the sampling scope to the most promising partition. The performance gap with other algorithms widens on sparse benchmarks and becomes greatest on Hartmann6-300D, where HiBO achieves over two magnitudes lower regrets than others. This indicates HiBO’s strong capacity to identify the most promising regions in high-dimensional tasks.
@@ 5.2.1 Evaluation Background
Besides evaluation on synthetic benchmarks, we evaluate the practical value of HiBO on a more complex real-world application, specifically automatic DBMS configuration tuning. The performance of DBMS highly depends on the settings of their configuration knobs, which control various aspects of their behaviours such as memory allocation, I/O behaviors, and query optimization. However, default configurations are typically suboptimal and require further tuning for optimal performance (Zhao et al., 2023). We select DBMS configuration tuning especially for the following considerations: 1) High-dimensionality. DBMS configuration tuning involves a high-dimensional search space with over 100 configuration knobs; 2) Knob Correlation. Many of these configuration parameters are interdependent, where certain knobs are correlated and can lead to combined effects on the performance; 3) Noisy Observations. Measurements of performance data in real-world environments are much noiser compared to cases in synthetic benchmarks. These factors make DBMS configuration tuning an ideal target for evaluating practical values of algorithms in real-world applications. In comparison, though not limited to simpler synthetic functions, commonly-seen “real-world" benchmarks in literature are still large based on simulation (Jones, 2008; Wang et al., 2017; Eriksson and Jankowiak, 2021). For comprehensiveness, we still included the evaluation on Mopta08 (Jones, 2008) and Rover Trajectory Planning (Wang et al., 2017) in Appendix E, including Mopta08 and Rover Trajectory Planning (Wang et al., 2017).
@@ 5.2.2 Experiment Setup
Our evaluation focuses on the effectiveness of HiBO for DBMS configuration tuning without relying on extra sampling besides its own basic initial sampling. We choose PostgreSQL as the target DBMS considering its accessibility and general applicability. From OnGres Inc. (2023), we manually selected 110 knobs for tuning, excluding those related to debugging, network connections, and authorization. SysBench (Kopytov, 2004) is applied as a standard OLTP workload in the experiment with throughput (measured in events per second, EPS) as the target metric to be maximized. We used a limited sample budget of 50 iterations for simulating scenarios in production environments.
@@ 5.2.3 Results and Analysis
Figure 4 (Left) summarizes the performance improvements of PostgreSQL on SysBench achieved by different algorithms. HiBO shows the highest throughput gains in greatest improvement rate, achieving about 28% of performance improvements over the default performance (200 EPS) than LA-MCTS and TuRBO do in 50 iterations. Moreover, without additional pre-sampling, OtterTune and LlamaTune also fall behind space-partitioning-based approaches. Such comparison highlights HiBO’s effectiveness even in complex and noisy real-world scenarios.
@@
Besides performance improvements, in real-world scenarios, tuning time and the rate of suggesting failure-causing configurations must be considered towards better practicality. A failure-causing configuration, defined as unsafe (Zhang et al., 2022), can degrade the performance of critical DBMS-dependent processes in the system. Therefore, we introduce the Safety-weighted Performance Improvement-Time Ratio (S-PITR): $S\text{-}PITR=\frac{PI}{TT+NF\cdot PE}$, where $PI$, $TT$, $NF$ and $PE$ refer to performance improvement, tuning time, number of failed (unsafe) configurations during the search and the penalty value for failed configurations respectively. This metric quantitatively measures the practicality of tuning algorithms by considering both tuning time and safety.
@@
Figure 4 (Right) presents the S-PITR values for the selected algorithms while tuning PostgreSQL’s configurations towards SysBench. Again, HiBO achieves the highest S-PITR value and outperforms LA-MCTS by over 20%. Besides, HiBO presents $\sim$25greater S-PITR scores than OtterTune and LlamaTune. These findings highlight HiBO’s adequate ability to strike a balance between performance improvement, safety, and efficiency in complex real-world DBMS tuning problems.
Figure 6: Comparison of HiBO and TuRBO across varying dimensionalities on the synthetic Levy benchmark. (Left) The ratio of mean regret (RMR) between TuRBO and HiBO ($\bar{regret}_{TuRBO}/\bar{regret}_{HiBO}$). (Right) The ratio of optimization time (ROT) for HiBO relative to TuRBO ($Time_{HiBO}/Time_{TuRBO}$).
5.3 Visual Analysis
@@
We visualize the acquisition function values generated by HiBO and TuRBO during the optimization of the Ackley2-200D synthetic function over 100 iterations. Starting with 20 initial samples, the distribution of acquisition values across the two effective dimensions as optimization progresses are plotted in Figure 5. For clarity, only the top 1000 acquisition values from 100x100 uniformly distributed samples are displayed at each iteration. The comparison shows that HiBO, leveraging space partitioning, converges its acquisition distribution more rapidly toward the optimal point, with the brightest areas near the optima. In contrast, TuRBO, while also moving toward the optimal region, converges more slowly and results in a less concentrated, noisier final distribution, with substantial acquisition values still scattered away from the optima. This analysis further validates the effectiveness of HiBO’s integration of space partitioning into the local optimizer’s acquisition strategy.
5.4 Empirical Validation on Scalability
To empirically evaluate HiBO’s scalability with increasing dimensionality, we compare the performance and timing costs of HiBO against plain TuRBO on the synthetic Levy benchmark across varying dimensions, as shown in Figure 6. Figure 6 (Left) depicts the ratio of mean regret (RMR) between TuRBO and HiBO ($\bar{regret}_{TuRBO}/\bar{regret}_{HiBO}$), which indicates HiBO’s performance advantage over TuRBO without the global-level navigator. It is observed that RMR increases steadily from 1.03 to 1.22 with dimensionality increasing from 50 to 1000, suggesting that TuRBO’s mean regret on Levy grows more significantly than HiBO’s with higher dimensionality, indicating that HiBO’s performance improvement scales effectively.
Figure 6 (Right) illustrates the ratio of optimization time (ROT) for HiBO relative to TuRBO ($Time_{HiBO}/Time_{TuRBO}$), reflecting the efficiency overhead brought by HiBO’s global-level navigator. Although ROT remains above 1 due to the additional time cost brought by the global-level navigator of HiBO, ROT decreases from 1.66 to 1.18 as dimensionality increasing. This indicates that as dimensionality grows, the local optimizer’s computational burden increases, while the global navigator’s overhead becomes less significant. As a result, the global navigator’s relative overhead is less significant as the local optimizer becomes the dominant computational component at higher dimensions.
@@ 6 Conclusion
In this paper, we introduce HiBO, a novel BO algorithm designed to efficiently handle high-dimensional search spaces via a hierarchical integration of global-level space partitioning and local modeling. Our comprehensive evaluation demonstrate its effectiveness and practical value on both synthetic benchmarks and real-world systems. Additionally, the visual analysis presents a straightforward validation of its effects in navigating on complex optimization landscapes. Its scalability w.r.t increased dimensionality is empirically validated on the synthetic benchmark. More discussions on the scalability of HiBO can be found in Appendix B.4.
@@ A.1 Algorithm Configurations for Synthetic Benchmarks
  • GP-BO: Implemented using ‘GaussianProcessRegressor‘ from scikit-learn library (Pedregosa et al., 2011), configured with a Matérn kernel ($\nu=2.5$) (Matérn, 2013) scaled by a constant factor of 1.0, and the noise level is set to 0.1, which translates to an alpha value of 0.01. The acquisition function is set to Expected Improvement (Jones et al., 1998);
@@ A.2 Experiment Setup for DBMS Configuration Tuning
We select PostgreSQL v16.3 (PostgreSQL Global Development Group, 2023) considering its accessibility and generality. PostgreSQL is a powerful, open-source Relational DBMS that has become one of the most popular choices for managing enterprise-level databases (Merit Data Tech, 2023) given its reliability and extensibility. Importantly, PostgreSQL exposes up to more than 200 configuration knobs as shown in PGCONF (OnGres Inc., 2023) which makes it an ideal instance for our DBMS configuration tuning experiments. We collected information of all configuration knobs of PostgreSQL in PostgreSQL Co.NF (OnGres Inc., 2023). A total of 110 knobs were manually selected, excluding those related to debugging, network connections, and authorization, as they could interfere with the interaction between the DBMS controller module and the DBMS instance and have negligible impact on DBMS performance.
@@ A.4 SysBench Statement Types
  • oint selection: "SELECT c FROM sbtest%u WHERE id=?"
@@ A.5 Hyperparameters
Appendix B Theoretical Analysis
B.1 Preliminaries and Assumption
Assumption B.1.
[…]
$\bigl{|}\mathbb{E}(\bar{v}-M_{v})\bigr{|}\;<\;\sigma_{v}.$ (4)
This assumption allows us to bound the fraction of solutions whose performance lies near the median or near the mean. It also induces an upper bound on how unevenly $\Omega$ can be split according to performance metrics.
B.2 Proof of Theorem 4.1
Proof.
Let $\delta_{1},\delta_{2},\ldots,\delta_{h}$ be the partition errors at successive tree levels along the path to the leftmost leaf. At the root, we split $\Omega$ into two sets of size at most $N/2+\delta_{1}$ and $N/2-\delta_{1}$, respectively. By assumption, the left child takes the larger set, hence it has size
$N/2\;+\;\delta_{1}.$
At the next level, we apply the same argument to that child’s sub-domain of size $N/2+\delta_{1}$, resulting in two further splits. In the worst case, the sub-domain with the highest partition score again takes the larger portion:
[…]
and so on, possibly incurring a partition error $\delta_{j}$ at each stage. Recursively unrolling this for $h$ splits, we get
$\underbrace{\delta_{h}\;+\;\frac{\delta_{h-1}}{2}\;+\;\frac{\delta_{h-2}}{2^{2}}\;+\;\dots\;+\;\frac{\delta_{1}}{2^{h-1}}}_{\text{accumulated partition error}}\;+\;\frac{N}{2^{h}}.$
Defining $\delta_{\max}=\max\{\delta_{1},\delta_{2},\dots,\delta_{h}\}$ and bounding each $\delta_{j}$ by $\delta_{\max}$ yields:
$\delta_{h}+\frac{\delta_{h-1}}{2}+\dots+\frac{\delta_{1}}{2^{h-1}}$ $\;\;\leq\;\;\delta_{\max}\,\Bigl{(}1\;\;+\dots\;\frac{1}{2^{\,h-1}}\Bigr{)}$
= $\;=\;2\,\delta_{\max}\,\Bigl{(}1-\frac{1}{2^{h}}\Bigr{)}.$
Combining this with $\tfrac{N}{2^{h}}$ completes the proof,
[…]
B.3 Comparison to LA-NAS and LA-MCTS
Single-Path vs. Weighted Sampling.
Methods like LA-NAS (Wang et al., 2021) and LA-MCTS (Wang et al., 2020) also partition the search space but only sample from the estimated most promising partition. For example, in LA-NAS’s theoretical analysis, all subsequent splits are localized to that single path, so that the domain of interest gets halved (plus partition error) at each level. This provides an exponential shrinkage bound (similar to Theorem 4.1) for that single partition.
HiBO’s Approach: Partition-Score Weighting.
By contrast, HiBO leverages the partition scores of all resulting leaf nodes for weighing the acquisition function values, instead of discarding sampling potential information from other partitions. This addresses a key limitation of single-path methods: if early splits are misleading, LA-NAS (or LA-MCTS) can end up discarding regions that later prove to contain globally better optima. In HiBO:
  • Leaves in which the local surrogate (and UCT) reveal good performance retain a high value of acquisition function.
  • Leaves with poor performance or high uncertainty gradually lose weight but are still revisited occasionally, mitigating the risk of “missing” hidden optima.
B.4 Scalability Analysis
An analysis about how the search-tree-based space partitioning scales with the dimensionality $d$ of the search space and the total number of samples $T$ is given below.
Dimension vs. Samples. For continuous or high-dimensional discrete $\Omega$, we typically let $N$ grow exponentially in $d$ (e.g., a grid with spacing $\Delta$ leads to grid $\text{grid size}\approx(1/\Delta)^{d}$). In such scenarios, any method that enumerates or partitions $\Omega$ naively could face exponential blow-up in $d$. HiBO mitigates this by:
  1. 1.
    Adaptive Expansion: Partitions are grown incrementally; subspaces with low potential (low UCT scores) are not expanded repeatedly.
  2. 2.
    Weighted Sampling: Unlike methods that only focus on a single path, HiBO keeps a global perspective to hedge against misleading early splits; yet high-performing partitions do receive exponentially more attention, so the search zooms in on promising regions quickly.
Consequently, while worst-case complexity may still be large for extremely high $d$, in practice HiBO often scales better than vanilla BO with full GP modeling, which suffers from $O(T^{3})$ kernel-inversion costs or from poor kernel fit in high dimensions.